# Spectral diversity in collisional neutrino-flavor conversion: flavor equipartition or swap

### v1

@@ I Introduction
On the other hand, the understanding of collisional flavor conversion (CFC) is still developing, except for CFI in the linear regime. Collisions with matter generally decohere neutrino correlation between two flavors, and suppression of FFI has been numerically observed [43, 44, 45, 46, 47, 48, 49, 50, 51]. However, when the flavor-dependent decoherence is coupled with each other through neutrino self-interactions, they mutate into a flavor instability [26]. From linear stability analysis, CFI has two distinct regimes called resonance-like and otherwise, which are also quite different in the growth rates [52, 53]. The growth rate of the standard CFI is dominated by collision rate $\propto R$ , whereas the resonance-like one is much faster and proportional to $\propto(\mu R)^{1/2}$, where $\mu$ denotes the self-interaction potential. Both of them have been searched in the state-of-the-art simulations of CCSNe and BNSM remnants, and potentially emerge at deep radii where neutrinos are less transparent against matter [54, 11, 52].
@@ II.1 Quantum kinetic equation
$\rho=\frac{1}{2}P_{0}\mathbbm{1}_{2}+\frac{f_{\nu_{e}}-f_{\nu_{x}}}{2}\bm{P}\cdot\bm{\sigma},$ (5)
where $\bm{\sigma}$ denotes the Pauli matrices, $\bm{P}$ the polarization vector, $P_{0}=\mathrm{Tr}\rho$, and $f_{\nu_{\alpha}}$ an occupation number distribution for a flavor $\alpha$. In our notations, the polarization vector is normalized by the difference between $\nu_{e}$ and $\nu_{x}$ in occupation distributions, and then the initial value of $P_{3}$, corresponding to the diagonal parts in the density matrix, is unity. Within the configuration, the QKE is recast into
@@ II.2 Linear stability analysis
$\Pi^{\mu\nu}(k)=\eta^{\mu\nu}+\sqrt{2}G_{\mathrm{F}}\int\mathrm{d}\Gamma_{\nu}^{\prime}\,(\rho_{ee}-\rho_{xx})\frac{v^{\mu}v^{\nu}}{v^{\sigma}k_{\sigma}+\mathrm{i}R_{E}}.$ (12)
@@
Eq. (13) can be analytically solved for a monochromatic neutrino distribution, and in the multi-energy case, the approximated solutions are provided with mean collision rates $\langle\overset{\scriptscriptstyle(-)}{R}\rangle$ in Refs. [55, 53]. The approximated solutions are
@@
$\omega_{\pm}\approx\begin{cases}-A-\mathrm{i}\gamma\pm\left(\left\lvert A\right\rvert+\mathrm{i}\dfrac{\left\lvert G\alpha\right\rvert}{\left\lvert A\right\rvert}\right)\,\,\,&\mathrm{if}\,A^{2}\gg\left\lvert G\alpha\right\rvert\\ -A-\mathrm{i}\gamma\pm\sqrt{\left\lvert G\alpha\right\rvert}\,\,\,&\mathrm{if}\,A^{2}\ll\left\lvert G\alpha\right\rvert.\end{cases}$ (16)
@@
$\mathrm{max}\left[\mathrm{Im}\,\omega_{\pm}\right]\approx-\gamma+\dfrac{\left\lvert G\alpha\right\rvert}{\left\lvert A\right\rvert}\,\,\,\mathrm{if}\,A^{2}\gg\left\lvert G\alpha\right\rvert,$ (17)
@@ II.3 Model
Figure 1 demonstrates collisional modes in flavor instability obtained numerically from Eq. (13) and solved analytically as Eq. (17) as a function of reaction rate $\bar{R}_{0}$ for antineutrinos. It is clearly found that dominant growing modes switch around $\langle\bar{R}_{0}\rangle\simeq\langle R_{0}\rangle$ as the reaction rate increases. As can be seen from the definition in Eqs. (15) and (16), the sign of $\alpha$ changes after that equality, and the maximum growing mode changes so as to compensate for that sign. Note that we fix neutrino number densities in this study, but if not so, the change of the sign of $A$ can bring similar behaviors [56]. The disparity in the sign between $A$ and $\alpha$ produces the distinction between the plus and minus modes. In the subsequent section, the two unstable modes display quite different behaviors in both linear and nonlinear dynamics.
@@ III Multi-Energy Flavor Evolution
Figure 3: Time evolution of the norm $\left\lvert\left\lvert\bm{P}\right\rvert\right\rvert$ of the polarization vector (top) and the angle of $P_{3}$ relative to the $z$-axis in the flavor space for neutrinos. In the bottom panel, all colored lines are overlapped.
@@
Since the flavor coherence $\left\lvert\left\lvert\bm{P_{\perp}}\right\rvert\right\rvert$ grows isoenergetically, the spectral structure of the norm shrinking is determined by the energy dependence of reaction rates $R_{E}$. So, the length of the polarization vector shortens monotonically with $\propto E_{\nu}^{2}$ in the energy domain. Meanwhile, the bottom panel of Fig. 3 exhibits the angle of $P_{3}$ relative to the third axis (hereafter denoted as the $z$-axis) in flavor space. It clearly shows that the tilt of the polarization vector is independent of neutrino energy and the asymptotic state for each neutrino is aligned upwardly along the $z$-axis. It means that leaning and standing the polarization vector are carried out by collisional decoherence shared through the self-interaction term. Actually, neutrinos with the lowest energy receive only the contribution from the self-interaction Hamiltonian $H_{\nu\nu}$, which explicitly does not depend on the neutrino spectra, because the collision rate is extremely small. Therefore, the asymptotic state is determined by the degree of collisional decoherence which they have experienced.
@@
$P_{ex}=\frac{1}{2}(P_{0}+P_{3})/P_{0}.$ (21)
The asymptotic states are now settled down in flavor eigenstates because the flavor coherence decays through collisional decoherence. Consequently, the norm of the polarization vector is almost identical to $P_{3}$ and determines the transition probability as follows:
$P_{ex}=\frac{1}{2}(1+\left\lvert\left\lvert\bm{P}\right\rvert\right\rvert).$ (22)
@@
However, linear and nonlinear behaviors significantly change when we employ lower antineutrino reaction rates. Figure 4 demonstrates the flavor evolution in the case of $\bar{R}_{0}=0.1\mathrm{\,km^{-1}}$. We find that neutrinos with the lowest energy establish a complete flavor swap, $P_{ex}=1$, while the higher energy one rather remains close to electron-type flavor, $P_{ex}<0.5$. This result is quite contrary to the case of $\bar{R}_{0}=1\mathrm{\,km^{-1}}$, and means that a flavor swap can occur even in the absence of the diagonal parts of collision term unlike the previous works [58] in the case with a resonance-like CFI. Also, in the bottom panel, the flavor coherence first undergoes linear damping isoenergetically, and then exponential growth with some energy spreading occurs. The behaviors can be understood through the unstable modes in the linear regime. As clarified in Fig. 1, this simulation setup has two types of non-stable modes, minus $\omega_{-}$ and plus modes $\omega_{+}$. As the antineutrino reaction rate decreases, the imaginary part $\mathrm{Im}(\omega_{-})$ of the minus mode reduces and becomes negative. On the other hand, that $\mathrm{Im}(\omega_{+})$ of the plus mode increases and switches into the growing mode. This means that the distinction between the two simulations above stems from the difference in the dominant growing modes.
@@ IV.1 Flavor Analysis by Eigenvector
The difference in the nonlinear behaviors between Figs. 2 and 4 can be understood by the spectral structure in the linear growth phase. In the case of $\bar{R}_{0}=1\mathrm{\,km^{-1}}$, all neutrinos with any energy simultaneously reach a linear saturation, and then collisional decoherence becomes to dominate the system evolution, as can be seen in Fig. 2. On the other hand, in the case of $\bar{R}_{0}=0.1\mathrm{\,km^{-1}}$, since the eigenvector of growing mode has a low-energy peaked distribution, neutrinos with lower energy reach the linear saturation before the higher energetic ones. The nonlinear power propagates into the other neutrinos via their self-interactions, and neutrinos with higher energy are forced to behave as in the nonlinear regime. However, the neutrinos themselves do not have enough significant flavor coherence $\rho_{ex}$, so the resultant collisional decoherence is still weak. As a consequence, collisional decoherence in the entire system can not prevent the low-energy neutrinos from undergoing a continuous flavor conversion in the nonlinear order, and they cannot establish a complete flavor swap.
@@
The balance between flavor coherence and collisional decoherence can be read out through the energy dependence of the norm collapse presented in Fig. 7. The distribution has the extremum in the middle of the considered neutrino energy range because the collision rate is proportional to $E_{\nu}^{2}$ whereas the flavor coherency is peaked at low energy. Actually, in the top panel of Fig.  5, neutrinos (thick green) with $\sim 20\mathrm{\,MeV}$ and antineutrinos (red) with $\sim 70\mathrm{\,MeV}$ are completely collapsed within the Bloch sphere. The fully depolarized (anti-)neutrinos establish a flavor equipartition as shown in Fig. 4. On the other hand, since neutrinos with the lowest energy continue to hold the length of the polarization vector, they can reach a full flavor swap. Also, neutrinos with the highest energy can not completely get away from shrinking the length of the polarization vectors, so they are at mixed states, which are not at pure states only with electron-type flavors.
@@ VI Acknowledgments
We thank Hiroki Nagakura, Lucas Johns, and Yudai Suwa for valuable comments and discussions. This work is supported by Grant-in-Aid for JSPS Fellows (Grant No. 22KJ2906) and JSPS KAKENHI Grant Number JP24H02245.

### v3

@@ I Introduction
On the other hand, the understanding of collisional flavor conversion (CFC) is still developing, except for CFI in the linear regime. Collisions with matter generally decohere neutrino correlation between two flavors, and suppression or damping of FFI has been numerically observed [43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. However, when the flavor-dependent decoherence is coupled with each other through neutrino self-interactions, they mutate into a flavor instability [26]. From linear stability analysis, CFI has two distinct regimes called resonance-like and otherwise, which are also quite different in the growth rates [53, 54]. The growth rate of the standard CFI is dominated by collision rate $\propto R$ , whereas the resonance-like one is much faster and proportional to $\propto(\mu R)^{1/2}$, where $\mu$ denotes the self-interaction potential. Both of them have been searched in the state-of-the-art simulations of CCSNe and BNSM remnants, and potentially emerge at deep radii where neutrinos are less transparent against matter [55, 11, 53].
@@ II.1 Quantum kinetic equation
$\rho=\frac{1}{2}P_{0}\mathbbm{1}_{2}+\frac{f_{\nu_{e}}^{i}-f_{\nu_{x}}^{i}}{2}\bm{P}\cdot\bm{\sigma},$ (5)
where $\bm{\sigma}$ denotes the Pauli matrices, $\bm{P}$ the polarization vector, $P_{0}=\mathrm{Tr}\rho$, and $f_{\nu_{\alpha}}$ an occupation number distribution for a flavor $\alpha$. In our notations, the polarization vector is normalized by the difference in the initial occupation number distributions between $\nu_{e}$ and $\nu_{x}$, and then the initial value of $P_{3}$, corresponding to the diagonal parts in the density matrix, is unity. Within the configuration, the QKE is recast into
@@ II.2 Linear stability analysis
$\Pi^{\mu\nu}(k)=\eta^{\mu\nu}+\sqrt{2}G_{\mathrm{F}}\int\mathrm{d}\Gamma_{\nu}\,(\rho_{ee}-\rho_{xx})\frac{v^{\mu}v^{\nu}}{v^{\sigma}k_{\sigma}+\mathrm{i}R_{E}}.$ (12)
@@
Eq. (13) can be analytically solved for a monochromatic neutrino distribution [26, 50], and in the multi-energy case, the approximated solutions are provided with mean collision rates $\langle\overset{\scriptscriptstyle(-)}{R}\rangle$ in Refs. [56, 54]. The approximated solutions are
@@
$\omega_{\pm}\approx\begin{cases}-A-\mathrm{i}\gamma\pm\left(\left\lvert A\right\rvert+\mathrm{i}\dfrac{G\alpha}{\left\lvert A\right\rvert}\right)\,\,\,&\mathrm{if}\,A^{2}\gg\left\lvert G\alpha\right\rvert\\ -A-\mathrm{i}\gamma\pm\sqrt{2\mathrm{i}G\alpha}\,\,\,&\mathrm{if}\,A^{2}\ll\left\lvert G\alpha\right\rvert.\end{cases}$ (16)
@@
$\mathrm{max}\left[\mathrm{Im}\,\omega_{\pm}\right]\approx-\gamma+\left\lvert\dfrac{G\alpha}{A}\right\rvert\,\,\,\mathrm{if}\,A^{2}\gg\left\lvert G\alpha\right\rvert,$ (17)
@@ II.3 Model
Figure 1 demonstrates collisional modes in flavor instability obtained numerically from Eq. (13) and solved analytically as Eq. (17) as a function of reaction rate $\bar{R}_{0}$ for antineutrinos. It is clearly found that dominant growing modes switch around $\langle\bar{R}_{E}\rangle\simeq\langle R_{E}\rangle$ with the reaction rate increasing. As can be seen from the definition in Eqs. (15) and (16), the sign of $\alpha$ changes after that equality, and the maximum growing mode changes so as to compensate for that sign. Note that we fix neutrino number densities in this study, but if not so, the change of the sign of $A$ can bring similar behaviors [57]. The disparity in the sign between $A$ and $\alpha$ produces the distinction between the plus and minus modes. In the subsequent section, the two unstable modes display quite different behaviors in both linear and nonlinear dynamics.
@@ III Multi-Energy Flavor Evolution
Figure 3: Time evolution of the norm $\left\lvert\left\lvert\bm{P}\right\rvert\right\rvert$ (top) and the angle $\theta_{3}$ (bottom) relative to the $z$-axis of the polarization vector $\bm{P}$ in the flavor space for neutrinos. In the bottom panel, all colored lines are overlapped.
@@
Since the flavor coherence $\left\lvert\left\lvert\bm{P_{\perp}}\right\rvert\right\rvert$ grows isoenergetically, the spectral structure of the norm shrinking is determined by the energy dependence of reaction rates $R_{E}$. So, the length of the polarization vector shortens monotonically with $\propto E_{\nu}^{2}$ in the energy domain. Meanwhile, the bottom panel of Fig. 3 exhibits the angle of the polarization vector relative to the third axis (hereafter denoted as the $z$-axis) in flavor space. It clearly shows that the tilt of the polarization vector is independent of neutrino energy and the asymptotic state for each neutrino is aligned upwardly along the $z$-axis. It means that leaning and standing the polarization vector are carried out by collisional decoherence shared through the self-interaction term. Actually, neutrinos with the lowest energy receive only the contribution from the self-interaction Hamiltonian $H_{\nu\nu}$, which explicitly does not depend on the neutrino spectra, because the collision rate is extremely small. Therefore, the asymptotic state is determined by the degree of collisional decoherence which they have experienced.
@@
$P_{ex}=\frac{1}{2}(P_{0}-P_{3})/P_{0}.$ (21)
The asymptotic states are now settled down in flavor eigenstates because the flavor coherence decays through collisional decoherence. Consequently, the length of the polarization vector is almost identical to $P_{3}$ and determines the transition probability as follows:
$P_{ex}=\frac{1}{2}(1-\left\lvert\left\lvert\bm{P}\right\rvert\right\rvert).$ (22)
@@
Figure 5: Energy distribution in the case of $\bar{R}_{0}=1$ (top) and $\bar{R}_{0}=0.1\,\mathrm{km^{-1}}$ (bottom). Dotted lines are initial spectra for electron-type neutrinos. Solid and dashed ones correspond to final one for neutrinos and antineutrinos after CFC, respectively. Clearly, the order of colored lines $\overset{\scriptscriptstyle(-)}{\nu}_{e}$ and $\overset{\scriptscriptstyle(-)}{\nu}_{x}$) is exchanged on the low-energy side between the two scenarios due to the presence or absence of collisional swap.
@@
However, linear and nonlinear behaviors significantly change when we employ lower antineutrino reaction rates. Figure 4 demonstrates the flavor evolution in the case of $\bar{R}_{0}=0.1\mathrm{\,km^{-1}}$. We find that neutrinos with the lowest energy establish a complete flavor swap, $P_{ex}=1$, while the higher energy one rather remains close to electron-type flavor, $P_{ex}<0.5$. This result is quite contrary to the case of $\bar{R}_{0}=1\mathrm{\,km^{-1}}$, and means that a flavor swap can occur even in the absence of the diagonal parts of collision term unlike the previous works [59] in the case with a resonance-like CFI. Also, in the bottom panel, the flavor coherence first undergoes linear damping isoenergetically, and then exponential growth with some energy spreading occurs. The behaviors can be understood through the unstable modes in the linear regime. As clarified in Fig. 1, this simulation setup has two types of non-stable modes, minus $\omega_{-}$ and plus modes $\omega_{+}$. As the antineutrino reaction rate decreases, the imaginary part $\mathrm{Im}(\omega_{-})$ of the minus mode reduces and becomes negative. On the other hand, that $\mathrm{Im}(\omega_{+})$ of the plus mode increases and switches into the growing mode. This means that the distinction between the two simulations above stems from the difference in the dominant growing modes. Comparing the energy distributions at the asymptotic states also clarifies the distinction as presented in Fig. 5. In the low-energy side, the order of final spectra of $\overset{\scriptscriptstyle(-)}{\nu}_{e}$ and $\overset{\scriptscriptstyle(-)}{\nu}_{x}$ are exchanged because collisional swap occurs in the case of $\bar{R}_{0}=0.1\,\mathrm{km^{-1}}$ (bottom). Also, in the case of $\bar{R}_{0}=1\,\mathrm{km^{-1}}$ (top), the complete overlap between two flavors appears in the high-energy side through a flavor equipartition.
@@ IV.1 Flavor Analysis by Eigenvector
The difference in the nonlinear behaviors between Figs. 2 and 4 can be understood by the spectral structure in the linear growth phase. In the case of $\bar{R}_{0}=1\mathrm{\,km^{-1}}$, all neutrinos with any energy simultaneously reach a linear saturation, and then collisional decoherence becomes to dominate the system evolution, as can be seen in Fig. 2. On the other hand, in the case of $\bar{R}_{0}=0.1\mathrm{\,km^{-1}}$, since the eigenvector of growing mode has a low-energy peaked distribution, neutrinos with lower energy reach the linear saturation before the higher energetic ones. The nonlinear power propagates into the other neutrinos via their self-interactions, and neutrinos with higher energy are forced to behave as in the nonlinear regime. However, the higher energetic neutrinos themselves do not have enough significant flavor coherence $\rho_{ex}$ yet, so the resultant collisional decoherence is still weak. As a consequence, collisional decoherence in the entire system can not prevent the low-energy neutrinos from undergoing a continuous flavor conversion in the nonlinear order, and they establish a complete flavor swap.
@@
The balance between flavor coherence and collisional decoherence can be read out through the energy dependence of the norm collapse presented in Fig. 8. The distribution has the extremum in the middle of the considered neutrino energy range because the collision rate is proportional to $E_{\nu}^{2}$ whereas the flavor coherency is peaked at low energy. Actually, in the top panel of Fig.  6, neutrinos (thick green) with $\sim 20\mathrm{\,MeV}$ and antineutrinos (red) with $\sim 70\mathrm{\,MeV}$ are completely collapsed within the Bloch sphere. The fully depolarized (anti-)neutrinos establish a flavor equipartition as shown in Fig. 4. On the other hand, since neutrinos with the lowest energy continue to hold the length of the polarization vector, they can reach a full flavor swap. Also, neutrinos with the highest energy can not completely get away from shrinking the length of the polarization vectors, so they are at mixed states, which are not at pure states only with electron-type flavors.
@@ VI Acknowledgments
We thank Hiroki Nagakura, Lucas Johns, and Yudai Suwa for valuable comments and discussions. This work is supported by Grant-in-Aid for JSPS Fellows (Grant No. 22KJ2906) and JSPS KAKENHI Grant Numbers JP24H02245 and JP25K17383.